OpenStudy (anonymous):

Lynne feeds her cat 65 g of cat food per day. About how many ounces of cat food will she feed her cat in 5 weeks? (1 oz ≈ 28.3 g)

OpenStudy (cloverracer):

First off, there are 7 days a week. How many days is that in 5 weeks?

OpenStudy (anonymous):

35

OpenStudy (anonymous):

hello

OpenStudy (cloverracer):

Good! Now how many times does 28.3 go into 65?

OpenStudy (anonymous):

29.9

OpenStudy (sjg13e):

\[?oz= 5weeks| \frac{ 7 days }{ 1 week }|\frac{ 65g }{ 1day }|\frac{ 1oz }{ 28.3g }=80.39oz \ in 5 weeks\]

OpenStudy (anonymous):

16.1 oz 80.4 oz 262.8 oz 1313.9 oz

OpenStudy (sjg13e):

Try the train track method. It makes conversions a lot easier. http://www.katmarsoftware.com/articles/railroad-track-unit-conversion.htm

OpenStudy (anonymous):

11.4

OpenStudy (cloverracer):

Please don't just give out answers, you have to make sure they understand. @realmadrid7

OpenStudy (anonymous):

im sorry m8

OpenStudy (anonymous):

@CloverRacer

OpenStudy (cloverracer):

It's okay